in a chess tournament, where the participants are to play one game with one another two player fell ill, having played only 3 games each .If the total numberof games played in the tournament is equal to 84 then find the number of participants in the beginning.     please tell me how it is 15

Grade:11

3 Answers

Riddhish Bhalodia
8 years ago
OK say there are x players from the start
thus x-2 non ill players
THe total matches they played with one another is
(x-2)(x-3)/2 …..........(because each player will play will every other player by (x-2)(x-3), but this will be double of the total answer so we half it)
in addition to this there are 6 more games played by ill players
\frac{(x-2)(x-3)}{2} + 6 = 84
this results in a quadratic as
x^2-5x -150 =0
whic results in
x=15
Ankit Jaiswal
8 years ago
x=15whic results inx^2-5x -150 =0this results in a quadratic as\frac{(x-2)(x-3)}{2} + 6 = 84hilet there are x players from the startthus x-2 non ill playersTHe total matches they played with one another is(x-2)(x-3)/2 …..........(because each player will play will every other player by (x-2)(x-3), but this will be double of the total answer so we half it)in addition to this there are 6 more games played by ill players
manmath
8 years ago
 
x=15 which results inx^2-5x -150 =0this results in a quadratic as\frac{(x-2)(x-3)}{2} + 6 = 84hilet there are x players from the startthus x-2 non ill playersTHe total matches they played with one another is(x-2)(x-3)/2 …..........(because each player will play will every other player by (x-2)(x-3), but this will be double of the total answer so we half it)in addition to this there are 6 more games played by ill players
 
2 days ago